### v2

# On Unitary Monodromy of Second-Order ODEs

Abstract
Given a second-order, holomorphic, linear differential equation $Lf=0$ on a Riemann surface, we say that its monodromy group $G\subset\operatorname{GL}(2,\mathbb{C})$ is unitary if it preserves a non-degenerate (though not necessarily positive) Hermitian form $H$ on $\mathbb{C}^{2}$ under the action $g\circ H\doteq g^{\dagger}Hg$. In the present work, we give two sets of necessary and sufficient conditions for a differential operator $L$ to have a unitary monodromy group, and we construct the form $H$ explicitly. First, in the case that the natural representation of $G$ on $\mathbb{C}^{2}$ is irreducible, we show that unitarity is equivalent to a set of easily-verified trace conditions on local monodromy matrices; in the case that it is reducible, we show that $G$ is unitary if and only if it is conjugate to a subgroup of one of two model subgroups of $\operatorname{GL}(2,\mathbb{C})$. Second, we show that unitarity of $G$ is equivalent to a criterion on the real dimension of the algebra $A$ generated by a rescaled group $G^{\prime}\subset\operatorname{SL}(2,\mathbb{C})$: that $\dim(A)=1$ if $G\subset S^{1}$ is scalar, $\dim(A)=2$ if $G$ is abelian, $\dim(A)=3$ if $G$ is non-abelian but its action on $\mathbb{C}^{2}$ is reducible, and $\dim(A)=4$ otherwise. We leverage these results to extend a conjecture of Frits Beukers on the spectrum of Lamé operators—namely, we give asymptotic and numerical evidence that his conjecture should apply similarly to a wider class of Heun operator. Our work makes progress towards characterizing the spectra of second-order operators on Riemann surfaces, and in particular, towards answering the accessory parameter problem for Heun equations.
I Introduction
Consider the equation $Lf=0$, where $L$ is a second-order, holomorphic differential operator on a (punctured) Riemann surface. We say the monodromy group $G\subset\operatorname{GL}(2,\mathbb{C})$ of this equation is unitary if it preserves a non-degenerate Hermitian form $H$ under the action $g:H\mapsto g^{\dagger}Hg$. We consider the following problem:
For which operators $L$ does the equation $Lf=0$ have a unitary monodromy group? (A)
In the present work, we give two sets of necessary and sufficient conditions for a general, second-order operator $L$ to have a unitary monodromy group, we construct the Hermitian form $H$ in both cases, and we demonstrate an application of our work to understand the point-spectra of differential operators. As a key example, we look at the Heun equation Heun (1888), which encompasses any second-order equation on $\mathbb{P}^{1}$ with at most four regular singularities. Heun’s equation takes the form
$\frac{d^{2}y}{dx^{2}}+\left(\frac{\gamma}{x}+\frac{\delta}{x-1}+\frac{\varepsilon}{x-a}\right)\frac{dy}{dx}+\frac{\alpha\beta x-\frac{B}{4}}{x(x-1)(x-a)}y=0,$ (1)
where $B\in\mathbb{C}$, and the parameters $\gamma,\delta,\varepsilon,\alpha,\beta\in\mathbb{R}$ are subject only to the restriction that $\gamma+\delta+\varepsilon=1+\alpha+\beta$. In this context, (1) can be reformulated as the accessory parameter problem:
For which values $B$ does (1) have a unitary monodromy group? (A’)
Even restricted to the Heun equation, this problem has several important applications. For one, the analytic Langlands correspondence claims that eigenfunctions of Hecke (or Baxter) and Gaudin operators on the moduli space of $\operatorname{PGL}_{2}$ bundles on the projective line, with parabolic structures at $N$ marked points, are parametrized by second order Fuchsian equations with appropriate singularities at these points and with real monodromy (i.e., taking values in $\operatorname{SL}_{2}(\mathbb{R})=\operatorname{SU}(1,1)$, which is unitary by Proposition II.2 below) (Etingof et al., 2024, Sec. 3.3). The simplest nontrivial example of this is given by $N=4$, which leads to the accessory parameter problem for the Heun (or Darboux) equation Etingof et al. (2021a, 2024). In a different direction, the accessory parameter problem for the Heun equation is equivalent to the problem of classifying spherical quadrilaterals. Indeed, each accessory parameter satisfying (A) corresponds to a unique spherical quadrilateral with ($\pi$-scaled) angles $(1-\gamma,1-\delta,1-\varepsilon,\alpha-\beta)$, up to congruence; Eremenko and Gabrielov used this correspondence to classify spherical rectangles, for which each angle has a half-integer value Eremenko and Gabrielov (2016).
Monodromy and the accessory parameter have also garnered wide interest in different contexts. Historically, monodromy of the Heun equation was studied in the context of isomonodromic deformations, tracing back to Paul Painlevé in ${}^{\text{th}}$ century. Specifically, Painlevé’s sixth equation characterizes the isomonodromic deformations of the Heun equation Takemura (2005). The accessory parameter has been studied as well—for instance, Nehari Nehari (1949) examined the impact of the accessory parameter on the Schwarz map generated by the ratio of two independent solutions to a Fuchsian equation. Additionally, Keen, Rauch, and Vasquez Keen et al. (1979) have studied the accessory parameter as a parameter relating different covering maps of the punctured torus, and Beukers Beukers (2002) has studied the accessory parameter’s impact on the $p$-adic radius of convergence for solutions to these differential equations.
A partial answer to (A) in the case of the Heun equation follows from classical work of Fricke and Klein (Fricke and Klein, 1897, p. 365–366). Namely, if $\gamma,\delta,\varepsilon,\alpha,\beta\in\mathbb{Q}$, then it is well-known that the monodromy group is determined by the so-called Fricke parameters; if $P,Q,R\in\operatorname{U}_{2}$ are the generators of the monodromy group, these are as follows:
$\operatorname{tr}(P),\operatorname{tr}(Q),\operatorname{tr}(R),\operatorname{tr}(PQ),\operatorname{tr}(QR),\operatorname{tr}(PR),\operatorname{tr}(PQR).$ (2)
In this context, if these seven values are known to be real and the monodromy representation is irreducible, one can deduce that the monodromy group itself is unitary (Reiter, 2011, Sec. 4). More generally, it was shown by Morgan and Shalen (Morgan and Shalen, 1984, Prop. III.1.1) and by Goldman (Goldman, 1988, Thm. 4.3) that a general subgroup $G\subset\operatorname{SL}(2,\mathbb{C})$ is unitary if and only if $\operatorname{tr}(g)\in\mathbb{R}$ for all $g\in G$, or equivalently, if the natural representation of $G$ on $\mathbb{C}^{2}$ has real character.
More recently, Beukers offered a stronger result in a special case of the Lamé equation Beukers (2007), which in turn is the case of (1) with $\gamma=\delta=\varepsilon=1/2$. Beukers showed that, if $\gamma=\delta=\varepsilon=1/2$ and $\alpha=\beta=1/4$, a unitary monodromy group is equivalent to a condition on only five of the seven Fricke parameters (2), and he leveraged this result to estimate the eigenvalues of the Lamé operator.
In the present work, we work towards answering (A) in the case of general second-order differential operators, by characterizing the unitary subgroups $G$ of $\operatorname{GL}(2,\mathbb{C})$. We provide two sets of equivalent criteria for unitarity. First, in the case where the natural representation of $G$ on $\mathbb{C}^{2}$ is irreducible, we show that unitarity is equivalent to a set of easily-verified trace conditions on local monodromy matrices, generalizing the result of Beukers on the Lamé equation and strengthening the result of Morgan and Shalen Morgan and Shalen (1984) and Goldman Goldman (1988) (by reducing the number of requisite trace conditions). In the case where the representation is reducible, we show that $G$ is unitary if and only if it is conjugate to a subset of one of two model subgroups of $\operatorname{GL}(2,\mathbb{C})$. Second, we show that unitarity is equivalent to a criterion on the real dimension of the algebra $A$ generated by a rescaled group $G^{\prime}\subset\operatorname{SL}(2,\mathbb{C})$: that $\dim(A)=1$ if $G\subset S^{1}$ is trivial, $\dim(A)=2$ if $G$ is abelian, $\dim(A)=3$ if $G$ is non-abelian but its representation on $\mathbb{C}^{2}$ is reducible, and $\dim(A)=4$ otherwise.
In all cases, we are able to explicitly construct the Hermitian form $H$ preserved by $G^{\prime}$ in terms of monodromy generators. We will see that a constructive proof is necessary to recover the second equivalent criterion (in terms of $\dim A$) above, and is instrumental in moving beyond the case of Heun operators. Connecting with another result of Beukers (Beukers, 2007, Prop. 4), quoted as Proposition IV.3 below, this also allows explicit construction of real-analytic solutions to second-order ODEs that admit them.
Finally, we apply our results to extend a conjecture of Beukers on the spectra of Lamé operators (Beukers, 2007, Conj. 1). Namely, leveraging a similar asymptotic estimate and numerical algorithm as Beukers, we provide evidence that the spectra of Heun operators should closely align with a (squared) parameter-dependent lattice. In doing so, we also provide strong evidence that our results can be strengthened in the case of Heun equations, with fewer trace conditions required to determine unitarity.
It has come to our attention during the preparation of this manuscript that Adachi Adachi (2022, 2023) has recently discovered an independent proof of our Theorem III.5.
II Background on Unitary Monodromy
We give a brief background on the central topics necessary for our analysis below. In particular, we briefly introduce the notions of monodromy groups and unitarity, and we outline how they apply to archetypal complex differential equations.
II.1 Monodromy and Local Exponents
Fix a compact Riemann surface $X$ and a finite set of punctures $\{z_{k}\}=S\subset X$, and consider the differential equation  [^1] Although $X$ compact and $S\subset X$ finite forms a primary case of interest, our results apply to general Riemann surfaces and any discrete collection of punctures.
$\frac{d^{2}y}{dz^{2}}+P(z)\frac{dy}{dz}+Q(z)y=0,$ (3)
where $P(z)$ and $Q(z)$ are holomorphic on $X\setminus S$. Let $\Gamma_{i}$ be closed, simple, counter-clockwise loops around each $z_{i}$ that do not encircle any $z_{j}\neq z_{i}$, each starting at a common point $p\notin S$.
Let $y_{1}$ and $y_{2}$ be two linearly independent solutions of (3) in a neighborhood $U\ni p$. We can analytically continue $y_{1}$ and $y_{2}$ about $\Gamma_{k}$, resulting in the new solutions $\widetilde{y}_{1}$ and $\widetilde{y}_{2}$ in $U$, respectively. In such a case, there exists an $M_{k}\in\operatorname{GL}(2,\mathbb{C})$ such that
$M_{k}\begin{pmatrix}y_{1}\\ y_{2}\end{pmatrix}=\begin{pmatrix}\widetilde{y}_{1}\\ \widetilde{y}_{2}\end{pmatrix}$
everywhere in $U$. The matrix $M_{k}$ is known as the monodromy matrix around $\Gamma_{k}$, and the group $G\subset\operatorname{GL}(2,\mathbb{C})$ generated by $\{M_{k}\}$ is known as the monodromy group of (3). Notably, the mapping
$\pi_{1}(X\setminus S)\to G,\;\Gamma_{k}\mapsto M_{k}$
is a group homomorphism.
A key application of monodromy matrices is that they quantify the asymptotic behavior of a differential equation near its singularities. In a sufficiently small neighborhood of a regular singularity $z_{k}$, we can generically find $y_{1}$ and $y_{2}$ such that $y_{1}=(z-z_{k})^{a_{k}}u_{1}$ and $y_{2}=(z-z_{k})^{b_{k}}u_{2}$, where $u_{1}$ and $u_{2}$ are nonzero analytic functions and $a_{k},b_{k}\in\mathbb{C}$. We refer to $a_{k}$ and $b_{k}$ as the local exponents of (3) at $z_{k}$. In general, if the monodromy matrix $M_{k}$ has eigenvalues $(e^{2\pi ia_{k}},e^{2\pi ib_{k}})$, we define the local exponents at $z_{k}$ (modulo $\mathbb{Z}$) to be $(a_{k},b_{k})$.
II.2 Unitarity
Given a nondegenerate Hermitian form $H$ on $\mathbb{C}^{2}$, we define the unitary group of $H$ to be
$\operatorname{U}(H)=\{g\in\operatorname{GL}(2,\mathbb{C})\;|\;g^{\dagger}Hg=H\},$
generalizing the classical unitary group $\operatorname{U}_{2}=\operatorname{U}(\mathbbm{1}_{2})$. In general, we define unitarity as follows:
Definition II.1.
A matrix group $G\subset\operatorname{GL}(2,\mathbb{C})$ is unitary if $G\subset\operatorname{U}(H)$ for a nondegenerate Hermitian form $H$.
Importantly, unitarity is preserved under conjugation; if $h\in\operatorname{GL}(2,\mathbb{C})$, then we have
$h^{-1}\operatorname{U}(H)h=\operatorname{U}(h^{\dagger}Hh).$
With this property in mind, we can break unitarity into two basic cases. If the form $H$ is (positive or negative) definite, it is conjugate to the Euclidean form ±$\pm\mathbbm{1}_{2}$; any group that preserves it must then be conjugate to a subgroup of $\operatorname{U}_{2}$. Similarly, if the form $H$ is indefinite, the unitary group $U(H)$ is given (up to conjugation) by the following well-known result (e.g., see Beukers (2007)):
Proposition II.2.
Let
$H_{0}=\begin{pmatrix}0&i\\ -i&0\\ \end{pmatrix}.$ (4)
Then $\operatorname{U}(H_{0})\simeq S^{1}\times\operatorname{SL}(2,\mathbb{R})$ is the group generated by $\operatorname{SL}(2,\mathbb{R})$ and the scalars $\lambda\in S^{1}$.
Remark.
Now, the standard embedding $S^{1}\simeq\operatorname{U}_{1}$ intersects $\operatorname{SL}(2,\mathbb{R})$ at the two points ±$\pm 1$; we clarify that the factor of $S^{1}$ in $\operatorname{U}(H_{0})\simeq S^{1}\times\operatorname{SL}(2,\mathbb{R})$ parameterizes the determinant of a matrix $g\in\operatorname{U}(H_{0})$, so that ±$\pm 1$ is in the kernel of the projection $\operatorname{U}(H_{0})\to S^{1}$.
II.3 The Equations of Lamé, Heun, and Darboux
Several aspects of our analysis are adapted from Frits Beukers’ work on the closely-related Lamé equation Beukers (2007). In algebraic form, Lamé’s equation reads as follows:
$p(x)\frac{d^{2}y}{dx^{2}}+\frac{1}{2}p^{\prime}(x)\frac{dy}{dx}+(Ax-B)y=0,$ (5)
for ‘accessory parameters’ $A,B\in\mathbb{C}$, and where $p(x)=4x(x-1)(x-a)$ and $x\in\mathbb{P}^{1}$. Comparing to (1), we see that Lamé’s equation is a special case of Heun’s equation with $\gamma=\delta=\varepsilon=1/2$; Beukers was interested in the particular subcase $A=1/4$, or equivalently, $\alpha=\beta=1/4$.
As with the Heun equation, the Lamé equation has four regular singularities: $z_{1}=0$, $z_{2}=1$, $z_{3}=a$, $z_{0}=\infty$. Unlike the Heun equation, however, the first three singularities all carry the same local exponents $(0,1/2)$. From the discussion above, this implies that the corresponding monodromy matrices $P=M_{1}$, $Q=M_{2}$, and $R=M_{3}$ have the eigenvalues ±$\pm 1$; we refer to such matrices as reflections below, though they may not be orthogonal reflections. Moreover, the product $M=PQR$ (which has local exponents $(1/4,1/4)$) is parabolic in the following sense:
Definition II.3.
A matrix $M$ is parabolic if it has two identical eigenvalues in $S^{1}$ and is not diagonalizable.
Using these properties of the monodromy group, Beukers was able to give a necessary and sufficient criterion for unitarity in the case of the Lamé equation with $A=1/4$; the following proposition is transcribed from (Beukers, 2007, Prop. 2):
Proposition II.4 (Beukers, 2007).
Suppose $P,Q,R\in\operatorname{GL}(2,\mathbb{C})$ are reflections, and their product $M=PQR$ is parabolic. Let $G$ be the group generated by $P$, $Q$, and $R$. Then the following statements are equivalent:
  1. 1.
    The group $G$ is unitary.
  2. 2.
    The traces $\operatorname{tr}(PQ)$, $\operatorname{tr}(QR)$, and $\operatorname{tr}(PR)$ are real.
  3. 3.
    The traces $t_{PQ}=\operatorname{tr}(PQ)$ and $t_{QR}=\operatorname{tr}(QR)$ are real, and satisfy $(t_{PQ}^{2}-4)(t_{QR}^{2}-4)\geq 16$.
In fact, this result provides significant insight into (A’) in Beukers’ case of interest. An accessory parameter $B$ gives rise to unitary monodromy if and only if the local monodromy matrices $P$, $Q$, and $R$ satisfy either of the above trace conditions. These trace conditions are easily verified on a computer, so Beukers was able to leverage his result to numerically estimate accessory parameters that give rise to unitary monodromy.
Recall that Heun’s equation (1) encompasses all second-order equations on $\mathbb{P}^{1}$ with at most four singular points, greatly generalizing the setting of Proposition II.4. With this level of generality, however, one sacrifices the strong constraints on local monodromy matrices that formed the basis of Beukers’ argument; instead, the local exponents of the Heun equation are given by (see, e.g., Xia et al. Xia et al. (2021))
At $\text{At }z=0,\quad\left(0,1-\gamma\right)$ (6)
At $\text{At }z=1,\quad\left(0,1-\delta\right)$
At $\text{At }z=a,\quad\left(0,1-\varepsilon\right)$
At $\text{At }z=\infty,\quad\left(\alpha,\beta\right).$
Although one might be interested in the general case $\gamma,\delta,\varepsilon,\alpha,\beta\notin\mathbb{R}$, these local exponents severely restrict the parameter space that could exhibit unitary monodromy:
Proposition II.5.
The monodromy group of the Heun equation (1) can only be unitary if $\gamma,\delta,\varepsilon\in\mathbb{R}$ and either $\alpha,\beta\in\mathbb{R}$ or $\alpha-\overline{\beta}\in\mathbb{Z}$.
Proof.
Write $P$, $Q$, and $R$ for the monodromy matrices about 00, $1$, and $a$, respectively, and suppose $G$ preserves a Hermitian form $H$. From the expressions (6), we find $\det(P)=e^{-2\pi i\gamma}$, and thus
$\det(P^{\dagger}HP)=|e^{-2\pi i\gamma}|^{2}\det(H)=e^{4\pi\operatorname{Im}(\gamma)}\det(H).$
Since $H$ must be preserved under the action of $G$, this implies that $\gamma\in\mathbb{R}$, and similarly for $\delta$, $\varepsilon$, and $\alpha+\beta$. It will follow from Lemma III.1 below that we must also have
$2\cos(\pi(\alpha-\beta))=\frac{\operatorname{tr}(PQR)}{\sqrt{\det(PQR)}}\in\mathbb{R},$
which can only be satisfied if either $\alpha-\beta\in\mathbb{R}$ or $\operatorname{Re}(\alpha-\beta)\in\mathbb{Z}$. These correspond to the two conditions given in the proposition statement. ∎
In the following section, we will see that this restriction falls out of our general characterization of unitary groups (Theorem III.5).
Finally, we note that the Heun equation is commonly expressed in terms of elliptic functions, using a change of variable discussed in Sirota and Smirnov (2006); Takemura (2010). For this, take values $\omega_{1},\omega_{2}\in\mathbb{C}$ (which we will fix shortly), and let $\wp(z)$ be the Weierstrass function Cartan (1995) with half-periods $\omega_{1}$, $\omega_{2}$, and $\omega_{3}=\omega_{1}+\omega_{2}$. Write $e_{i}=\wp(\omega_{i})$, and fix $\omega_{i}$ such that
$a=\frac{e_{2}-e_{1}}{e_{3}-e_{1}}.$
Substitute $\wp(z)=e_{1}+(e_{2}-e_{1})x$, and let
$\begin{gathered}m_{0}=\alpha-\beta-\frac{1}{2},\qquad m_{1}=\frac{1}{2}-\gamma,\\ m_{2}=\frac{1}{2}-\delta,\qquad m_{3}=\frac{1}{2}-\varepsilon.\end{gathered}$ (7)
Finally, write
$u=x^{-m_{1}/2}(x-1)^{-m_{2}/2}(x-a)^{-m_{3}/2}y.$
With this change of variables, we convert (1) into the Darboux equation:
$\frac{d^{2}u}{dz^{2}}-\left(\sum_{i=0}^{3}m_{i}(m_{i}+1)\wp(z-\omega_{i})\right)u=B^{\prime}u,$ (8)
where we write $\omega_{0}=0$. Here, $B^{\prime}=(e_{3}-e_{1})B+B_{0}$ for a fixed $B_{0}=B_{0}(\alpha,\beta,\gamma,\delta,\varepsilon)$. A full expression of $B_{0}$ is given in Takemura (2004), for instance, but we do not make use of it here. The Darboux equation has the local exponents
At $\text{At }z=\omega_{i},\quad(-m_{i},1+m_{i}),$ (9)
which can be deduced from (6) by applying the described coordinate transform. Notably, all four monodromy matrices have unit determinant, so the proof of Proposition (II.5) cannot be applied here; indeed, we will see in Section (IV) that Darboux’s equation satisfies (A) for a far wider range of parameters than Heun’s equation.
The Lamé equation (5) takes on a particularly simple form in these coordinates, as $m_{1}$, $m_{2}$, and $m_{3}$ all vanish:
$\frac{d^{2}u}{dz^{2}}+A\wp(z)u=B^{\prime}u,$ (10)
noting that $u(z)=y(x)$ in this case.
III Unitary Subgroups of $\operatorname{GL}(2,\mathbb{C})$
GL(2,C) We can make progress on the problem (A) by considering general matrix groups. Namely, the monodromy group of any second-order equation is a subgroup $G\subset\operatorname{GL}(2,\mathbb{C})$, up to conjugacy class, so an answer to the following question provides a criterion for any second-order operator to satisfy (A):
Which subgroups $G\subset\operatorname{GL}(2,\mathbb{C})$ are unitary, in the sense of Definition II.1? (B)
We answer this question completely in the present section. Proposition III.3 first answers the question for groups $G\subset\operatorname{SL}(2,\mathbb{C})$ generated by three elements, as we encounter in the Heun (1) and Darboux (8) equations; this case follows from the general results of Morgan and Shalen (Morgan and Shalen, 1984, Prop. III.1.1) and Goldman (Goldman, 1988, Thm. 4.3), but we will see that the constructive proof we offer of Proposition III.3 is necessary to extend to arbitrary subgroups of $\operatorname{GL}(2,\mathbb{C})$. Theorem III.5 carries out exactly extension. In short, for any subgroup $G\subset\operatorname{GL}(2,\mathbb{C})$, we define a rescaled subgroup
$G^{\prime}=\left\{\pm g/\sqrt{\det g}\;\big{|}\;g\in G\right\}\subset\operatorname{SL}(2,\mathbb{C}).$
Notably, $G$ is unitary if and only if $G^{\prime}$ is unitary and $|\det g|=1$ for all $g\in G$, so (B) can be answered for subgroups $G^{\prime}\subset\operatorname{SL}(2,\mathbb{C})$ without any loss of generality.
In one direction, we will show that for subgroups $G\subset\operatorname{GL}(2,\mathbb{C})$ for which the natural representation of $G$ on $\mathbb{C}^{2}$ is reducible, the rescaled group $G^{\prime}$ must be contained within either $\operatorname{SO}_{2}$ or the group ×$\mathbb{R}\rtimes\mathbb{R}^{\times}\subset\operatorname{SL}(2,\mathbb{R})$ of real, upper-triangular matrices; if the representation of $G$ on $\mathbb{C}^{2}$ is irreducible, we will show that unitarity is instead equivalent to a set of easily-verified trace conditions on an arbitrary set $S\subset G$ of generators. In a different direction, we show that unitarity is equivalent to a strict criterion on the real dimension of the algebra $A$ generated by $G^{\prime}$. In both cases, we identify the Hermitian form $H$ preserved by $G$.
Our results provide insight into the point-spectra of any second-order holomorphic differential operator; we study two archetypal examples in the following section. For the Lamé equation (5), where $G$ is generated by a set $\{P,Q,R\}$ of complex reflections whose product $PQR$ is parabolic, our result corresponds to that of Beukers Beukers (2007), written above as Proposition II.4. For the Heun equation (1) with $\alpha,\beta,\gamma,\delta,\varepsilon\in\mathbb{Q}$ and a monodromy group $G$ acting irreducibly on $\mathbb{C}^{2}$, our result corresponds to the criterion for unitarity using Fricke parameters (Reiter, 2011, Sec. 4).
One direction of our general theorem follows straightforwardly from Proposition II.2:
Lemma III.1.
Suppose $G\subset\operatorname{SL}(2,\mathbb{C})$ is unitary, in the sense of Definition II.1. Then all elements of $G$ have real trace.
Proof.
Suppose $G\subset\operatorname{U}(H)$ for a nondegenerate Hermitian form $H$. If $H$ is indefinite, for one, it is conjugate to the form $H_{0}$ of Proposition II.2. But then $G$ is conjugate to a subgroup of $\operatorname{U}(H_{0})\simeq\operatorname{U}_{1}\otimes\operatorname{SL}(2,\mathbb{R})$; since $G\subset\operatorname{SL}(2,\mathbb{C})$ by hypothesis, it must in fact be conjugate to a subgroup of $\operatorname{SL}(2,\mathbb{R})$, from which the claim follows.
If $H$ is definite, it is conjugate to the Euclidean form (up to sign), and $G$ is conjugate to a subgroup of $\operatorname{SU}_{2}$. Any element of $\operatorname{SU}_{2}$ has eigenvalues $\lambda,\lambda^{-1}\in S^{1}$, which implies our result. ∎
The converse of Lemma III.1 does not generally hold: indeed, consider the matrices
$P=\begin{pmatrix}1&z_{p}\\ &1\end{pmatrix},\qquad Q=\begin{pmatrix}1&z_{q}\\ &1\end{pmatrix},\qquad z_{p}\overline{z}_{q}\notin\mathbb{R}.$
It is clear that $P$, $Q$, and all products thereof have real trace. However, suppose they both preserve a Hermitian form
$H=\begin{pmatrix}h_{11}&h_{12}\\ \overline{h}_{12}&h_{22}\end{pmatrix},$
so that $P^{\dagger}H=HP^{-1}$ and $Q^{\dagger}H=HQ^{-1}$. Matching elements shows that $h_{11}=0$, implying $h_{12}\neq 0$, as well as $\overline{z}_{p}h_{12}=-z_{p}\overline{h}_{12}$ and $\overline{z}_{q}h_{12}=-z_{q}\overline{h}_{12}$. These conditions can only be met if $z_{p}\overline{z}_{q}\in\mathbb{R}$, and we reach a contradiction.
Over the next few lemmas, we demonstrate that, by and large, this is the only way for the converse of Lemma III.1 to fail. As a first step, if we restrict ourselves to three-element-generated groups acting irreducibly on $\mathbb{C}^{2}$, the converse holds. As discussed above, the next two results—Lemma III.2 and Proposition III.3—follow from the work of Morgan and Shalen (Morgan and Shalen, 1984, Prop. III.1.1) and Goldman (Goldman, 1988, Thm. 4.3). Even still, in order to generalize to Theorem III.5 below, we require a constructive proof of these statements; that is, given matrices $P,Q,R\in\operatorname{SL}(2,\mathbb{C})$, we must identify the particular Hermitian form $H$ they preserve. We carry out this program below:
Lemma III.2.
Fix matrices $P,Q,R\in\operatorname{SL}(2,\mathbb{C})$ with $\operatorname{tr}(P)\neq\pm 2$, and let $G$ be the group generated by $\{P,Q,R\}$. Suppose the natural representation of $G$ on $\mathbb{C}^{2}$ is irreducible. Then $G$ is unitary if and only if
$\operatorname{tr}(P),\operatorname{tr}(Q),\operatorname{tr}(R),\operatorname{tr}(PQ),\operatorname{tr}(QR),\operatorname{tr}(PR),\operatorname{tr}(PQR)\in\mathbb{R}.$ (11)
Proof.
[…]
Since $\det(P)=1$ and $\operatorname{tr}(P)\neq\pm 2$, the matrix $P$ is necessarily diagonalizable. By conjugating our matrices appropriately, we can put $P$ in the form
$P=\begin{pmatrix}\lambda_{P}&\\ &\lambda_{P}^{-1}\end{pmatrix},\qquad\lambda_{P}\neq\pm 1,$ (12)
and in the same basis, we write
$Q=\begin{pmatrix}q_{11}&q_{12}\\ q_{21}&q_{22}\end{pmatrix},\qquad R=\begin{pmatrix}r_{11}&r_{12}\\ r_{21}&r_{22}\end{pmatrix}.$ (13)
We split the remainder of our proof into cases, depending on whether $|\operatorname{tr}(P)|>2$ or $|\operatorname{tr}(P)|<2$.
First case. Suppose that $|\operatorname{tr}(P)|>2$, so that, in particular, $\lambda_{P}\in\mathbb{R}$. Since we know that $\lambda_{P}\neq\pm 1$, the conditions on $\operatorname{tr}(Q)$ and $\operatorname{tr}(PQ)$ imply that $q_{11},q_{22}\in\mathbb{R}$, and similarly with $r_{11},r_{22}\in\mathbb{R}$. But then, as
$\begin{gathered}\operatorname{tr}(QR)=q_{11}r_{11}+q_{22}r_{22}+q_{12}r_{21}+q_{21}r_{12}\in\mathbb{R},\\ \operatorname{tr}(PQR)=\lambda_{P}q_{11}r_{11}+\lambda_{P}^{-1}q_{22}r_{22}+\lambda_{P}q_{12}r_{21}+\lambda_{P}^{-1}q_{21}r_{12}\in\mathbb{R},\end{gathered}$ (14)
we also see that $q_{12}r_{21},q_{21}r_{12}\in\mathbb{R}$. Now, consider a conjugation of our group by the matrix
$T=\begin{pmatrix}t^{-\frac{1}{2}}&\\ &t^{\frac{1}{2}}\end{pmatrix},$ (15)
with $t\in\mathbb{C}$. This transformation preserves $P$ and the diagonal elements of $Q$ and $R$, but maps $q_{12}\mapsto q_{12}t^{-1}$ and $q_{21}\mapsto q_{21}t$, and likewise for $R$; notably, it leaves $q_{12}r_{21}$ and $q_{21}r_{12}$ unchanged.
If we first suppose that $q_{12},q_{21}\neq 0$, we can fix a value $t$ such that $q_{12}\in\mathbb{R}$, and thus that $r_{21}\in\mathbb{R}$. Since $\det(Q)=1\in\mathbb{R}$ and both diagonal elements of $Q$ are real, we must also have $q_{21}\in\mathbb{R}$ (and thus $r_{12}\in\mathbb{R}$) in this basis. A similar argument holds if any of the pairs $\{r_{12},r_{21}\}$, $\{q_{21},r_{12}\}$, or $\{q_{12},r_{21}\}$ are nonzero, or if at most one of $\{q_{12},q_{21},r_{12},r_{21}\}$ is nonzero.
The final possibility is that either $q_{12},r_{12}\neq 0$ or $q_{21},r_{21}\neq 0$, but the remaining off-diagonal elements vanish. This is prevented by the hypothesis that the representation of $G$ on $\mathbb{C}^{2}$ is irreducible.
Second case. Suppose now that $|\operatorname{tr}(P)|<2$, so that $\lambda_{P}\in S^{1}$. We check that $G$ preserves a diagonal Hermitian form
$H=\begin{pmatrix}h_{11}&\\ &h_{22}\end{pmatrix},$
for as-of-yet unfixed (but nonzero) values $h_{11},h_{22}\in\mathbb{R}$. We expand out the equation $Q^{\dagger}H\overset{?}{=}HQ^{-1}$ as follows:
$\begin{pmatrix}h_{11}\overline{q}_{11}&h_{22}\overline{q}_{21}\\ h_{11}\overline{q}_{12}&h_{22}\overline{q}_{22}\end{pmatrix}\overset{?}{=}\begin{pmatrix}h_{11}q_{22}&-h_{11}q_{12}\\ -h_{22}q_{21}&h_{22}q_{11}\end{pmatrix},$ (16)
and likewise for $R$. Matching matrix elements, we see that $Q,R\in\operatorname{U}(H)$ if and only if
$q_{11}=\overline{q}_{22},\qquad r_{11}=\overline{r}_{22},\qquad q_{12}=-h_{22}\overline{q}_{21}/h_{11},\qquad r_{12}=-h_{22}\overline{r}_{21}/h_{11}.$
Now, the conditions on $\operatorname{tr}(Q)$ and $\operatorname{tr}(PQ)$ imply that $q_{11}=\overline{q}_{22}$, and similarly, $r_{11}=\overline{r}_{22}$. As such, the conditions (14) show that $q_{21}r_{12}=\overline{q_{12}r_{21}}$.
If none of the off-diagonal elements vanish, we fix $h_{11}=1$ and
$h_{22}=-\frac{r_{12}}{\overline{r}_{21}}=-\frac{q_{12}}{\overline{q}_{21}}.$
Indeed, this value must be real, as
$r_{12}r_{21}=r_{11}r_{22}-1\in\mathbb{R}.$
Suppose instead that $q_{12},q_{21}\neq 0$, but either $r_{12}=0$ or $r_{21}=0$; the identity $q_{21}r_{12}=\overline{q_{12}r_{21}}$ then shows that $R$ is diagonal; the choice $h_{22}=-q_{12}/\overline{q}_{21}$ works as before. A similar argument holds in the case $r_{12},r_{21}\neq 0$, and the same identity prevents the case $q_{12},r_{21}\neq 0$, $q_{21},r_{12}=0$ or its reverse. Finally, the irreducibility hypothesis prevents either $q_{21}=r_{21}=0$ or $q_{12}=r_{12}=0$. ∎
We now extend our result to the case where $\operatorname{tr}(P)=\pm 2$; since $P$ may no longer be diagonalizable, we apply somewhat different techniques to recover a preserved Hermitian form.
Proposition III.3.
Fix matrices $P,Q,R\in\operatorname{SL}(2,\mathbb{C})$, and let $G$ be the group generated by $\{P,Q,R\}$. Suppose the representation of $G$ on $\mathbb{C}^{2}$ is irreducible. Then $G$ is unitary if and only if
$\operatorname{tr}(P),\operatorname{tr}(Q),\operatorname{tr}(R),\operatorname{tr}(PQ),\operatorname{tr}(QR),\operatorname{tr}(PR),\operatorname{tr}(PQR)\in\mathbb{R}.$
Proof.
If any of $M\in\{P,Q,R\}$ satisfy $\operatorname{tr}(M)\neq\pm 2$, we can cycle the matrices until $\operatorname{tr}(P)\neq\pm 2$, and apply Lemma III.2; indeed, note that the trace conditions  (11) are closed under cyclic permutations. As such, we suppose without loss of generality that $\operatorname{tr}(P),\operatorname{tr}(Q),\operatorname{tr}(R)=2$; otherwise, we replace matrices with their negations as appropriate (which changes neither the trace hypothesis nor the unitarity condition).
Suppose now that $\operatorname{tr}(PQ)\neq\pm 2$, and define the matrices
$P^{\prime}=PQ,\qquad Q^{\prime}=Q^{-1},\qquad R^{\prime}=R.$ (17)
These matrices generate the group $G$, and they satisfy
$\operatorname{tr}(P^{\prime}),\operatorname{tr}(Q^{\prime}),\operatorname{tr}(R^{\prime}),\operatorname{tr}(P^{\prime}Q^{\prime}),\operatorname{tr}(P^{\prime}R^{\prime}),\operatorname{tr}(P^{\prime}Q^{\prime}R^{\prime})\in\mathbb{R}$
by the conditions (11), noting that $\operatorname{tr}(Q^{-1})=\operatorname{tr}(Q)=2$. It remains to check $\operatorname{tr}(Q^{\prime}R^{\prime})$. For this, recall that every matrix satisfies its own characteristic polynomial—in particular,
$Q^{2}-2Q+1=0,$
or equivalently,
$Q+Q^{-1}=2.$
Then we find
$\operatorname{tr}(Q^{\prime}R^{\prime})=\operatorname{tr}(Q^{-1}R)=2\operatorname{tr}(R)-\operatorname{tr}(QR)\in\mathbb{R}.$
As such, the matrices (17) satisfy the hypotheses of Lemma III.2, proving unitarity. A similar argument applies to $QR$, $PR$, and $PQR=P^{\prime}R^{\prime}$, so we can assume that
$\operatorname{tr}(PQ),\operatorname{tr}(QR),\operatorname{tr}(PR),\operatorname{tr}(PQR)=\pm 2.$
Suppose also that $P$ is parabolic, in the sense of Definition II.3; if none of $\{P,Q,R\}$ are parabolic, the result follows trivially (but also, the irreducibility hypothesis is violated). Bring $P$ to Jordan normal form as
$P=\begin{pmatrix}1&1\\ &1\end{pmatrix},$ (18)
and write $Q$ and $R$ as in (13).
We suppose first that $\operatorname{tr}(PQ)=\operatorname{tr}(PR)=\operatorname{tr}(QR)=-2$ and $\operatorname{tr}(PQR)=2$, so that
$q_{21}=\operatorname{tr}(PQ)-\operatorname{tr}(Q)=-4,\qquad r_{21}=\operatorname{tr}(PR)-\operatorname{tr}(R)=-4=q_{21},$
and consequently,
$\operatorname{tr}(QR)=q_{11}r_{11}+q_{22}r_{22}+q_{12}r_{21}+q_{21}r_{12}$ $=q_{11}r_{11}+q_{22}r_{22}+q_{12}q_{21}+r_{12}r_{21}$
$=q_{11}r_{11}+q_{22}r_{22}+q_{11}q_{22}+r_{11}r_{22}-2$
$=(q_{11}+r_{22})(q_{22}+r_{11})-2=-2,$
applying $\det(Q)=\det(R)=1$. This implies either $q_{11}+r_{22}=0$ or $q_{22}+r_{11}=0$, but we also see that
$\operatorname{tr}(PQR)-\operatorname{tr}(QR)=q_{22}r_{21}+q_{21}r_{11}=-4(q_{22}+r_{11})=4,$
so that $q_{22}+r_{11}=-1$. From above, then, $q_{11}+r_{22}=0$, so we reach a contradiction:
$14=\operatorname{tr}(Q)+\operatorname{tr}(R)=q_{11}+q_{22}+r_{11}+r_{22}=-$ 1 .
As such, we see that either $\operatorname{tr}(PQR)=-2$ or one of $\operatorname{tr}(PQ)$, $\operatorname{tr}(PR)$, or $\operatorname{tr}(QR)$ is $2$. In the former case, we make the replacement
$P^{\prime\prime}=-PQ,\qquad Q^{\prime\prime}=Q^{-1},\qquad R^{\prime\prime}=R.$
These matrices satisfy $\operatorname{tr}(P^{\prime\prime})=\operatorname{tr}(Q^{\prime\prime})=\operatorname{tr}(R^{\prime\prime})=2$, by hypothesis, and
$\operatorname{tr}(P^{\prime\prime}Q^{\prime\prime})=-\operatorname{tr}(PQR)=2.$
In either case, then, we can suppose that
$\operatorname{tr}(P)=\operatorname{tr}(Q)=\operatorname{tr}(R)=\operatorname{tr}(PQ)=2.$
The same analysis as above shows that $q_{21}=0$ in this basis, and thus (from the condition $\det(Q)=1$) that $q_{11}=q_{22}=1$. Since $\operatorname{tr}(PR)=2$ would similarly imply that $r_{21}=0$, and thus violate our irreducibility hypothesis, we take $\operatorname{tr}(PR)=-2$ and thus $r_{21}=-4$.
In this case, we have $\operatorname{tr}(QR)=\pm 2$, and so
$\operatorname{tr}(QR)-\operatorname{tr}(R)=q_{12}r_{21}=-4q_{12}=\pm 2-2,$
implying either $q_{12}=1$ (and so $Q=P$) or $q_{12}=0$ (and so $Q=1$). In either case, the only conditions on $R$ are that $\operatorname{tr}(R)=2$, $\det(R)=1$, and $\operatorname{tr}(PR)=-2$, so it can take the general form
$R=\begin{pmatrix}1+2s&s^{2}\\ -4&1-2s\end{pmatrix},\qquad s\in\mathbb{C}.$
All three of $\{P,Q,R\}$ then preserve the Hermitian form
$H=\begin{pmatrix}\operatorname{Im}(s)&-i\\ i&\end{pmatrix},$
proving the theorem. ∎
We move on now to the case where the representation of $G$ on $\mathbb{C}^{2}$ is reducible, first showing that any unitary group must be conjugate to a subgroup of one of two model groups:
Lemma III.4.
Suppose $G\subset\operatorname{SL}(2,\mathbb{C})$, and suppose the natural representation of $G$ on $\mathbb{C}^{2}$ is reducible. In this case, $G$ is unitary if and only if it is conjugate to a subgroup of either $\operatorname{SO}(2,\mathbb{R})\subset\operatorname{SL}(2,\mathbb{C})$ or the group ×$\mathbb{R}\rtimes\mathbb{R}^{\times}\subset\operatorname{SL}(2,\mathbb{C})$ of real upper-triangular matrices.
Proof.
First, suppose $G$ is unitary, and fix $P,Q\in G$. By the reducibility hypothesis, we can conjugate $G$ so that all $g\in G$ are upper-triangular. If $\operatorname{tr}(P)\neq\pm 2$, we can bring it into the form (12) while retaining the upper-triangular form of $Q$.
Suppose $G$ preserves a Hermitian form
$H=\begin{pmatrix}h_{11}&h_{12}\\ \overline{h}_{12}&h_{22}\end{pmatrix},$ (19)
so that $P^{\dagger}H=HP^{-1}$. If $|\operatorname{tr}(P)|>2$, matching elements shows that $h_{11}=h_{22}=0$; in particular, we can conjugate $G$ by the matrix (15) to bring $H$ to the form (4), or a real multiple thereof. Proposition II.2 then shows that $G$ is conjugate to a subset of the group ×$\mathbb{R}\rtimes\mathbb{R}^{\times}\subset\operatorname{SL}(2,\mathbb{C})$ of real upper-triangular matrices.
If $|\operatorname{tr}(P)|<2$, matching elements of $P^{\dagger}H=HP^{-1}$ instead shows that $h_{12}=0$, and subsequently matching elements of $Q^{\dagger}H=HQ^{-1}$ shows that $Q$ must be diagonal:
$Q=\begin{pmatrix}\lambda_{Q}&\\ &\lambda_{Q}^{-1}\end{pmatrix}.$
But from Lemma III.1, we know that $\operatorname{tr}(P)$, $\operatorname{tr}(Q)$, and $\operatorname{tr}(PQ)$ must be real; since $|\operatorname{tr}(P)|<2$, we have $\lambda_{P}\in S^{1}$, and these conditions show that $\lambda_{Q}\in S^{1}$ as well. Of course, the group of matrices of this form is simply an eigendecomposition of $\operatorname{SO}(2,\mathbb{R})$.
The alternative is that $\operatorname{tr}(P)=\pm 2$ for every $P\in G$. Fix $P,Q\in G$ such that $P\neq\pm 1$; if this is not possible, then $G$ is trivial. Noting that both of our model groups are closed under negation, suppose without loss of generality that $\operatorname{tr}(P)=\operatorname{tr}(Q)=2$, so that
$P=\begin{pmatrix}1&z_{p}\\ &1\end{pmatrix},\qquad Q=\begin{pmatrix}1&z_{q}\\ &1\end{pmatrix},\qquad z_{p},z_{q}\in\mathbb{C}.$
Moreover, conjugating by the matrix (15) allows us to fix $z_{p}=1$. Suppose that $P$ and $Q$ preserve the form (19). Matching elements yields $h_{11}=0$, implying $h_{12}\neq 0$, as well as $h_{12}=-\overline{h}_{12}$ and $\overline{z}_{q}h_{12}=-z_{q}\overline{h}_{12}$; these conditions require $z_{q}\in\mathbb{R}$, which implies our result. The converse—that our two model groups are unitary—follows from Proposition II.2. ∎
Finally, we treat general subgroups of $\operatorname{GL}(2,\mathbb{C})$. In the reducible case, we use the classification of Lemma III.4 to deduce the dimensionality of $\mathbb{R}\otimes G$ as a real algebra; since conjugation is an isomorphism of algebras, this property is basis-independent. To extend the irreducible case beyond the hypotheses of Proposition III.3, we make use of the fact that its proof is constructive; given $P$, $Q$, and $R$, we can deduce the Hermitian form $H$ that they preserve. Indeed, even two of these matrices determine $H$ uniquely, so we can study a generic $R^{\prime}\in G$ by looking at the group generated by $\{P,Q,R^{\prime}\}$. We prove the following result; for convenience, we say that $G$ is reducible if the natural representation of $G$ on $\mathbb{C}^{2}$ is reducible, and irreducible otherwise.
Theorem III.5.
Suppose $G\subset\operatorname{GL}(2,\mathbb{C})$, and define the rescaled group
$G^{\prime}=\left\{\pm g/\sqrt{\det g}\;\big{|}\;g\in G\right\}.$ (20)
Let $A$ be the real algebra generated by $G^{\prime}$; if $G\subset\operatorname{SL}(2,\mathbb{C})$ already, we can equivalently consider $G^{\prime}=G$. Then $G$ is unitary if and only if $|\det g|=1$ for all $g\in G$ and one of the following (mutually exclusive) criteria is satisfied:
  1. 1.
    $G\subset S^{1}$ is scalar; equivalently, $\dim(A)=1$.
  2. 2.
    $G$ is abelian (and thus reducible), and $\dim(A)=2$. Equivalently, $G^{\prime}$ is conjugate to a non-scalar subgroup of $\operatorname{SO}(2,\mathbb{R})$, of the group ×$\mathbb{R}^{\times}\subset\operatorname{SL}(2,\mathbb{C})$ of real diagonal matrices, or of the group $\mathbb{Z}_{2}\times\mathbb{R}\subset\operatorname{SL}(2,\mathbb{C})$ of real upper-triangular matrices of trace ±$\pm 2$.
  3. 3.
    $G$ is reducible but non-abelian, and $\dim(A)=3$. Equivalently, $G^{\prime}$ is conjugate to a non-abelian subgroup of the group +×$\mathbb{R}^{+}\rtimes\mathbb{R}^{\times}\subset\operatorname{SL}(2,\mathbb{C})$ of real upper-triangular matrices.
  4. 4.
    $G$ is irreducible, and $\dim(A)=4$. Equivalently, $G$ is conjugate to an irreducible subgroup of either $\operatorname{SL}(2,\mathbb{R})$ or of $\operatorname{U}_{2}$. Equivalently, $G^{\prime}$ is irreducible, and, for a generating set $S\subset G^{\prime}$ and all $P,Q,R\in S$, we have
    $\operatorname{tr}(P),\operatorname{tr}(Q),\operatorname{tr}(R),\operatorname{tr}(PQ),\operatorname{tr}(PR),\operatorname{tr}(QR)\in\mathbb{R}$
    and either $\operatorname{tr}(PQR)\in\mathbb{R}$ or $\operatorname{tr}(PRQ)\in\mathbb{R}$.
Remark.
We will see that, if $G$ is irreducible and we can identify non-commuting elements $P,Q\in S$ with $|\operatorname{tr}(P)|\geq 2$, we can take $P$ and $Q$ to be fixed in the above trace conditions.
Proof.
If any $g\in G$ has $|\det g|\neq 1$, then the argument of Lemma II.5 shows that unitarity is impossible; we thus assume $|\det g|=1$ for all $g\in G$. Furthermore, if $G$ is abelian, reducible, or unitary, then the rescaled group $G^{\prime}$ shares the same property, and conversely. As such, we suppose without loss of generality that $G\subset\operatorname{SL}(2,\mathbb{C})$.
Suppose first that $G$ is unitary, reducible, and abelian, but contains a non-scalar element; by Lemma III.4, we know that $G$ is (up to conjugation) contained either in $\operatorname{SO}_{2}\subset\operatorname{SL}(2,\mathbb{C})$, in the subgroup $\mathbb{Z}_{2}\times\mathbb{R}\subset\operatorname{SL}(2,\mathbb{C})$ of real upper-triangular matrices of trace ±$\pm 2$, or in the subgroup ×$\mathbb{R}^{\times}\subset\operatorname{SL}(2,\mathbb{C})$ of real diagonal matrices. All three generate an algebra of real dimension $2$. Similarly, if $G$ is unitary, reducible, and non-abelian, then $G$ is contained (up to conjugation) in the group ×$\mathbb{R}\rtimes\mathbb{R}^{\times}\subset\operatorname{SL}(2,\mathbb{C})$ of real upper-triangular matrices, which generates an algebra of real dimension $3$. If we fix non-commuting $P,Q\in G$, then the algebra generated by $\{1,P,Q\}\subset G$ is of real dimension $\geq 3$, and the claim follows.
Conversely, suppose $G$ is reducible and abelian and that $\dim(A)=2$. Fix $P\in G$ with $P\neq\pm 1$; if this is not possible, then $G$ is scalar. Now, if $\operatorname{tr}(P)\notin\mathbb{R}$, then $P+P^{-1}=2\operatorname{tr}(P)$ is linearly independent from $1$ over the reals; diagonalizing $P$, we thus see that $A$ contains the ($4$ dimensional) algebra of complex diagonal matrices. As such, we know that $\operatorname{tr}(P)\in\mathbb{R}$. If $\operatorname{tr}(P)\neq\pm 2$, we can write $P$ in the form (12); otherwise, we can write $P$ in the form (18), up to sign. Now, the set $\{1,P\}$ already generates an algebra of real dimension $2$, in both cases. Indeed, if $|\operatorname{tr}(P)|>2$, this set spans the algebra of real diagonal matrices; if $|\operatorname{tr}(P)|<2$, it spans the same algebra as $\operatorname{SO}_{2}$; and if $|\operatorname{tr}(P)|=2$, it spans the algebra of real upper-triangular matrices with repeated eigenvalues. By our hypothesis, then, $A$ (and thus $G$) is contained in this span. But Theorem III.5 shows that the intersection of any of these three algebras with $\operatorname{SL}(2,\mathbb{C})$ is unitary, so the same must be true of $G$.
Next, suppose that $G$ is reducible and non-abelian and that $\dim(A)=3$. Fix non-commuting $P,Q\in G$; following our argument above, we know that $\operatorname{tr}(P),\operatorname{tr}(Q),\operatorname{tr}(PQ)\in\mathbb{R}$. If $\operatorname{tr}(P)\neq\pm 2$, we again write $P$ in the form (12)); writing $Q$ as (13), the reducibility hypothesis allows us to fix $q_{21}=0$. If $|\operatorname{tr}(P)|>2$, the derived trace conditions show that the diagonals of both $P$ and $Q$ are real; if $|\operatorname{tr}(P)|<2$, they show that all diagonal elements are contained in $S^{1}$. In either case, the non-commuting condition requires $q_{12}\neq 0$, so, by conjugating $G$ appropriately, we fix $q_{12}=1$. As such, the real algebra $A^{\prime}$ generated by $\{1,P,Q\}$ contains the space of real, strictly upper-triangular matrices. If any of the diagonal elements of $P$ or $Q$ are not real, then $A^{\prime}$ must also contain the space of complex, strictly upper-triangular matrices, and we see that $\dim(A^{\prime})\geq 4$. Otherwise, $A^{\prime}$ is the algebra of real upper-triangular matrices (of real dimension $3$), and the claim follows. Finally, suppose without loss of generality that $\operatorname{tr}(P)=2$, and write $P$ in the form (18); as before, write $Q$ as (13) with $q_{21}=0$. For $P$ and $Q$ not to commute, then, we must have $q_{11},q_{22}\neq\pm 1$. If $q_{11},q_{22}\notin\mathbb{R}$, the same argument as before demonstrates that the real algebra $A^{\prime}$ generated by $\{1,P,Q\}$ has $\dim(A^{\prime})\geq 4$, and we reach a contradiction. Thus, $q_{11},q_{22}\in\mathbb{R}$, and the only way to satisfy $\dim(A^{\prime})=3$ is to have $q_{12}\in\mathbb{R}$ as well.
[…]
Suppose that $|\operatorname{tr}(P)|\neq 2$, and conjugate $G$ to bring $P$ to the form (12); as before, write $Q$ as (13). If $|\operatorname{tr}(P)|>2$, then the trace conditions on $P$ and $Q$ guarantee that $\lambda_{P},q_{11},q_{22}\in\mathbb{R}$, and, with an additional conjugation by the matrix (15), we can further ensure that $q_{12},q_{21}\in\mathbb{R}$. This is clear if one of the two elements vanishes—note that they cannot both vanish, by the condition $PQ\neq QP$. If neither element vanishes, we can choose $q_{12}\in\mathbb{R}$ and apply the condition $\det(Q)=1$ to ensure $q_{21}\in\mathbb{R}$. Without loss of generality, we suppose $q_{12}\neq 0$.
Choose an $R\in S$, and note that $r_{11},r_{22}\in\mathbb{R}$. Since $q_{12}\neq 0$, then the conditions on $\operatorname{tr}(QR)$ and $\operatorname{tr}(PQR)$ (or $\operatorname{tr}(PRQ)$) show that $q_{12}r_{21},q_{21}r_{12}\in\mathbb{R}$, and thus $r_{21}\in\mathbb{R}$. If $r_{21}\neq 0$, the condition $\det(R)=1$ thus implies that $r_{12}\in\mathbb{R}$. Otherwise, fix a matrix
$M=\begin{pmatrix}m_{11}&m_{12}\\ m_{21}&m_{22}\end{pmatrix}\in S$
with $m_{21}\neq 0$; the existence of such a matrix follows from our irreducibility hypothesis. As before, we find that $m_{11},m_{22},m_{12},m_{21}\in\mathbb{R}$, so our conditions on $\operatorname{tr}(MR)$ and $\operatorname{tr}(PMR)$ (or $\operatorname{tr}(PRM)$) imply that $r_{12}\in\mathbb{R}$. Since $R$ is general, and our conjugation depends only on the matrices $P$ and $Q$, we find $G\subset\operatorname{SL}(2,\mathbb{R})$, so Proposition II.2 shows that $G$ is unitary.
If $|\operatorname{tr}(P)|<2$, the same trace conditions show that $\lambda_{P}\in S^{1}$ and $q_{11}=\overline{q}_{22}$. Choose an $R\in S$; then the trace conditions now show that $r_{11}=\overline{r}_{22}$ and $q_{21}r_{12}=\overline{q_{12}r_{21}}$. If $q_{12}$ (resp., $q_{21}$) vanishes, we see that $r_{12}$ (resp., $r_{21}$) must vanish as well; but $R$ was generic, so this would violate our irreducibility hypothesis. Thus, $q_{12},q_{21}\neq 0$, and the set $\{P,Q,R\}$ preserves the form (19) with $h_{21}=0$, $h_{11}=1$, and $h_{22}=-q_{12}/\overline{q}_{21}\in\mathbb{R}$. Since both our conjugation into this basis and the form $H$ depend only on $P$ and $Q$, our result follows.
The final case is, whenever $P,Q\in S$ are non-commuting, we have $\operatorname{tr}(P),\operatorname{tr}(Q)=\pm 2$; without loss of generality, suppose $\operatorname{tr}(P)=\operatorname{tr}(Q)=2$. By hypothesis, $P,Q\neq 1$, so we can bring $P$ to the form (18). In this basis, we must have $q_{21}\neq 0$; otherwise, the conditions $\det(Q)=1$ and $\operatorname{tr}(Q)=2$ would imply $q_{11}=q_{22}=1$, which violates the non-commuting assumption.
Now, $P$ also cannot commute with $PQ$, so that $\operatorname{tr}(PQ)=\pm 2$. As in Proposition III.3, the non-commuting condition requires $\operatorname{tr}(PQ)=-2$, and thus $q_{21}=-4$; then $Q$ takes the following form:
$Q=\begin{pmatrix}1+2s&s^{2}\\ -4&1-2s\end{pmatrix},\qquad s\in\mathbb{C}.$
Now, $P$ and $Q$ uniquely preserve the Hermitian form (19) with $h_{11}=\operatorname{Im}(s)$, $h_{12}=-i$, and $h_{22}=0$. However, Proposition III.3 shows that, for any $R\in S$, the group generated by $\{P,Q,R\}$ is unitary. Thus, $R$ preserves the same Hermitian form, and $G$ is unitary.
Finally, we prove that, if $G$ is irreducible, then it is unitary if and only if $\dim(A)=4$. Suppose that $G$ is irreducible and unitary; by an appropriate conjugation, we can ensure either that $G\subset\operatorname{SL}(2,\mathbb{R})$ or that $G\subset\operatorname{SU}_{2}$, and so $\dim(A)\leq 4$. By the density theorem Etingof et al. (2011), however, we know that $G$ must generate the full matrix algebra $\mathbb{C}^{2\times 2}$ over the complex numbers; since this algebra is of (complex) dimension 4, the real algebra $A$ must be of real dimension $\geq 4$. Alternatively, we can derive this result more directly. In the case that $G\subset\operatorname{SL}(2,\mathbb{R})$, the only way for $\dim(A)<4$ is for all $P\in G$ to be upper-triangular (or lower-triangular), which would imply that $G$ is reducible. In the case that $G\subset\operatorname{SU}_{2}$, fix non-commuting $P,Q\in G$, and conjugate by a unitary matrix to ensure $P$ takes the form (12); that this is possible follows from the spectral theorem. Since $q_{11}=\overline{q}_{22}$, we can subtract off the diagonal of both $Q$ and $PQ$ using an appropriate linear combination of $\{1,P\}$; but $\lambda_{P}\neq\pm 1$, so the algebra $A^{\prime}$ generated by $\{1,P,Q\}$ must contain all matrices of the following form:
$Z=\begin{pmatrix}z&w\\ -\overline{w}&\overline{z}\end{pmatrix}.$
Since $A^{\prime}\subset A$, this implies that $\dim(A)=4$. We note that, in this latter case, $A$ is isomorphic to the quaternion algebra.
Conversely, suppose that $G$ is irreducible and that $\dim(A)=4$. If any $P\in G$ satisfies $\operatorname{tr}(P)\notin\mathbb{R}$, then $P+P^{-1}=2\operatorname{tr}(P)$ is linearly independent from $1$ over the reals; but this implies that the algebra $A^{\prime}$ generated by $\{1,P\}$ already has $\dim(A^{\prime})=4$, so $A^{\prime}=A$. But then, any $Q\in G$ must be contained in this algebra, so, in particular, $G$ must be reducible. As such, we know that $\operatorname{tr}(P)\in\mathbb{R}$ for all $P\in G$, which implies our result. ∎
IV Unitary Monodromy of Heun and Darboux Operators
Although we have approached the problem (A) in the general case with Theorem III.5, it is worth discussing how our result applies to the Heun (1) and Darboux (8) operators specifically. In doing so, we will be able to make progress towards characterizing the point-spectra of Heun and Darboux operators.
For one, since the monodromy matrices of the Darboux equation are all of unit determinant, we can apply Theorem III.5 with $G^{\prime}=G$:
Corollary IV.1.
Let $G$ be the monodromy group of the Darboux equation (8). If the natural representation of $G$ on $\mathbb{C}^{2}$ is reducible, then $G$ is unitary if and only if it is conjugate to a subgroup of either $\operatorname{SO}(2,\mathbb{R})$ or the group ×$\mathbb{R}\rtimes\mathbb{R}^{\times}\subset\operatorname{SL}(2,\mathbb{C})$ of real upper-triangular matrices.
If the representation of $G$ on $\mathbb{C}^{2}$ is irreducible, fix local monodromy matrices $P,Q,R\in G$. Then $G$ is unitary if and only if the following conditions are met:
  1. 1.
    For each $i\in\{0,1,2,3\}$, either $m_{i}\in\mathbb{R}$ or $\operatorname{Re}(m_{i})\in\frac{1}{2}\mathbb{Z}$.
  2. 2.
    The traces $\operatorname{tr}(PQ)$, $\operatorname{tr}(QR)$, and $\operatorname{tr}(PR)$ are real, or equivalently, the matrices $\{P,Q,R\}$ generate a real algebra of dimension $4$.
Proof.
That the remaining trace conditions of Theorem III.5 are equivalent to our conditions on the parameters $m_{i}$ follows from our expressions (9) for local exponents. ∎
Before applying the lemma to the Heun equation (1), we note that—except for particular choices of parameters—the representation of $G$ on $\mathbb{C}^{2}$ cannot be reducible. Indeed, suppose it is reducible, and suppose $u\in\mathbb{C}^{2}$ is a shared eigenvector of all $g\in G$. Fix local monodromy matrices $P$, $Q$, and $R$ about $z=0$, $1$, and $a$, respectively. On one hand, the expressions (6) show that the eigenvalue of $PQR$ associated with $u$ must be $e^{2\pi i\alpha}$, without loss of generality. On the other, we know that the eigenvalue of $P$ associated with $u$ must be either $1$ or $e^{2\pi i\gamma}$, and similarly with $Q$ and $R$; multiplying these matrices together, we see that the representation can only be reducible if
$\alpha\in\{0,\gamma,\delta,\varepsilon,\gamma+\delta,\gamma+\varepsilon,\delta+\varepsilon,\gamma+\delta+\varepsilon\}\pmod{1}.$ (21)
In particular, this criterion ensures that the representation is irreducible in the Lamé case (5) with $A=1/4$, where $\gamma=\delta=\varepsilon=1/2$ and $\alpha=\beta=1/4$; in this case, our argument reduces to one of Beukers Beukers (2007). It similarly ensures irreducibility if $\alpha,\beta\notin\mathbb{R}$. In any case, we can apply Theorem III.5 once again:
Corollary IV.2.
Let $G$ be the monodromy group of the Heun equation (1), and suppose (21) is not satisfied. Fix local monodromy matrices $P,Q,R\in G$ and define $P_{0}=e^{\pi i\gamma}P$, $Q_{0}=e^{\pi i\delta}Q$, and $R_{0}=e^{\pi i\varepsilon}R$. Then $G$ is unitary if and only if the following conditions are met:
  1. 1.
    The parameters $\gamma$, $\delta$, and $\varepsilon$ are real, and either $\alpha,\beta\in\mathbb{R}$ or $\alpha-\overline{\beta}\in\mathbb{Z}$.
  2. 2.
    The traces $\operatorname{tr}(P_{0}Q_{0})$, $\operatorname{tr}(Q_{0}R_{0})$, and $\operatorname{tr}(P_{0}R_{0})$ are real, or equivalently, the matrices $P_{0}$, $Q_{0}$, and $R_{0}$ generate a real algebra of dimension 4.
As discussed previously, this result encompasses both Proposition II.5 and Beukers’ result on the Lamé equation Beukers (2007), quoted as Proposition II.4 above. The former follows from the first criterion of Theorem III.5, that is, that the rescaled group $G^{\prime}$ associated to the monodromy group $G$ is itself a subgroup of $\operatorname{SL}(2,\mathbb{C})$. The latter result follows from fixing the parameter values $\alpha=\beta=1/4$, $\gamma=\delta=\varepsilon=1/2$, and applying the trace identity (Beukers, 2007, Lemma 2).
It is worth noting that the setting of Corollary IV.1 is somewhat more general than that of Corollary IV.2. In particular, the Heun parameters $\gamma,\delta,\varepsilon\in\mathbb{R}$ correspond to Darboux parameters $m_{1},m_{2},m_{3}\in\mathbb{R}$, as we can see from the transformations (7), and the restrictions on $\alpha$ and $\beta$ correspond to the parameter space $m_{0}\in\mathbb{R}\cup\left(\frac{1}{2}+\mathbb{Z}+i\mathbb{R}\right)$. By contrast, our result on the Darboux equation applies to the larger space $m_{i}\in\mathbb{R}\cup\left(\frac{1}{2}\mathbb{Z}+i\mathbb{R}\right)$ for all parameters.
To make sense of this, recall that the variable $z$ of (8) and the variable $x$ of (1) are related as $x\sim z^{2}$ about every pole. If $G_{H}$ is the monodromy group of a fixed Heun equation and $G_{D}$ is that of the corresponding Darboux equation, then so long as both act irreducibly on $\mathbb{C}^{2}$, we see that the resulting map $G_{D}\hookrightarrow G_{H}$ is a strict inclusion; a simple closed loop in $G_{D}$ corresponds to two loops in $G_{H}$. As a result, if $G_{H}$ is unitary, we expect the same of $G_{D}$; if only $G_{D}$ is unitary, however, we cannot generally expect the same of $G_{H}$.
We can make more progress on this question by characterizing the point-spectrum (i.e., set of eigenvalues) of Heun and Darboux operators. To this end, we transcribe Proposition IV.3 below from (Beukers, 2007, Prop. 4):
Proposition IV.3.
Let $G$ be the monodromy group of the linear second order differential equation $y^{\prime\prime}+py^{\prime}+qy=0$, where $p,q\in\mathbb{C}(z)$. Then $G$ is unitary if and only if there exists a nonzero $C^{2}$ function $f$ on $\mathbb{C}\setminus\{0,1,a\}$ which is a real-analytic solution of the equation. Furthermore, if the natural representation of $G$ on $\mathbb{C}^{2}$ is irreducible, then $f$ is uniquely determined up to a constant factor.
Remark.
In fact, Beukers’ proof holds for the case where $p$ and $q$ are simply analytic on our (punctured) domain, so it applies equally well to the Darboux equation (8).
With this result in hand, we see that Theorem III.5 provides a criterion for any $\lambda\in\mathbb{C}$ to belong to the point-spectrum of any second-order operator. In a different direction, however, it allows us to connect the Heun and Darboux operators through the coordinate transform of Section II.3.
Naïvely, the change of variables does not take one real-analytic function to another, as it maps
$y(x)\mapsto x^{-m_{1}/2}(x-1)^{-m_{2}/2}(x-a)^{-m_{3}/2}y(x)=:u$
for real values $m_{1}$, $m_{2}$, and $m_{3}$; however, we can correct this behavior by instead mapping
$y(x)\mapsto|x|^{-m_{1}}|x-1|^{-m_{2}}|x-a|^{-m_{3}}y(x),$ (22)
which differs from the above by only an antiholomorphic factor (and thus continues to satisfy the Darboux equation). As such, Proposition IV.3 yields the following corollary:
Corollary IV.4.
If a Heun equation (1) carries a unitary monodromy group with irreducible action on $\mathbb{C}^{2}$, then the corresponding Darboux equation (8)—resulting from the transform outlined in Section II.3—also carries a unitary monodromy group.
The converse does not hold, as discussed above: not every real-analytic solution of the Darboux equation corresponds to a real-analytic solution for the Heun equation. In the change of variables taken to reach the Darboux equation, note that $z$ and $-z$ have the same pre-image $x=(\wp(z)-e_{1})/(e_{2}-e_{1})$. In moving from the Heun to the Darboux equation, then, we introduce an additional requirement that the solution is even—in general, solutions of the Heun equation (1) correspond to even solutions of the Darboux equation (8).
V Estimating the Eigenvalues of Heun Operators
Following Beukers Beukers (2007), we provide an asymptotic estimate for the eigenvalues of the Heun equation, and we compare this estimate to a numerical approximation of the same eigenvalues. In so doing, we argue that Beukers’ derived estimate should extend to any second-order equation with at most four regular singularities.
As a starting point, we convert the Heun equation (1) into the Darboux equation (8) through the change of variable defined in Section II.3; this procedure yields a lattice $\Lambda\subset\mathbb{C}$ with half-periods $\omega_{1}$, $\omega_{2}$, and $\omega_{3}=\omega_{1}+\omega_{2}$, and parameters $m_{0}\in\mathbb{R}\cup(\frac{1}{2}+\mathbb{Z}+i\mathbb{R})$ and $m_{1},m_{2},m_{3}\in\mathbb{R}$. Notably, these parameters satisfy the first criterion of Corollary IV.1, so we might expect to find unitary monodromy—or equivalently, single-valued real-analytic solutions (Proposition IV.3).
In this setting, we extend and revise Beukers’ conjecture on the distribution of eigenvalues $B$ of Heun’s operator. Importantly, recall from Section II.3 that these values are distinct from the eigenvalues $B^{\prime}$ of Darboux’s operator; if such a $B$ exists, then a corresponding $B^{\prime}$ exists (Corollary IV.4) and is given by $B^{\prime}=(\wp(\omega_{3})-\wp(\omega_{1}))B+B_{0}$, where $B_{0}=B_{0}(\alpha,\beta,\gamma,\delta,\varepsilon)$. These values $B^{\prime}$ correspond to single-valued, real-analytic, even solutions of the Darboux equation.
We split the modified conjecture into two distinct claims. First, we note that the argument underlying (Beukers, 2007, Conj. 1) can be straightforwardly adapted to support the following conjecture:
Conjecture V.1.
Let $\overline{\Lambda}$ be the lattice generated by $2\overline{\omega_{1}}$ and $2\overline{\omega_{2}}$ (i.e., the conjugate of the lattice $\Lambda$ defined above), and let
$\Delta\doteq 4\operatorname{Im}(\overline{\omega}_{1}\omega_{2})=2i(\omega_{1}\overline{\omega}_{2}-\omega_{2}\overline{\omega}_{1})$
be the area of its fundamental parallelogram. Furthermore, let $\zeta$ be the Weierstrass zeta function Cartan (1995) on $\Lambda$, and let $\eta_{i}=\zeta(z+2\omega_{i})-\zeta(z)$ be its quasi-periods. Let $\ell_{0}\in\frac{2\pi}{\Delta}\overline{\Lambda}$, and consider the limit $\ell_{0}\to\infty$. Fix an accessory parameter $B$ such that Heun’s equation carries a single-valued, real-analytic solution, and let $B^{\prime}=(\wp(\omega_{3})-\wp(\omega_{1}))B+B_{0}$ be the corresponding accessory parameter for the Darboux equation (8). If $B^{\prime}$ satisfies $\sqrt{B^{\prime}}=\ell_{0}+\varepsilon$ for some $\varepsilon\in(-\pi/\Delta,\pi/\Delta]\times(-\pi/\Delta,\pi/\Delta]$, then it must be given by
$B^{\prime}=\ell_{0}^{2}-\frac{2}{i\Delta}\left(\sum_{i=0}^{3}m_{i}(m_{i}+1)\right)\left(\eta_{1}\overline{\omega_{2}}-\eta_{2}\overline{\omega_{1}}+\pi i\frac{\ell_{0}}{\overline{\ell_{0}}}\right)+\mathcal{O}(1/|\ell_{0}|),$
or equivalently (if $\ell_{0}\neq 0$),
$\sqrt{B^{\prime}}=\ell_{0}-\frac{1}{i\Delta\ell_{0}}\left(\sum_{i=0}^{3}m_{i}(m_{i}+1)\right)\left(\eta_{1}\overline{\omega_{2}}-\eta_{2}\overline{\omega_{1}}+\pi i\frac{\ell_{0}}{\overline{\ell_{0}}}\right)+\mathcal{O}(1/|\ell_{0}|^{2}).$
Remark.
This estimate is a special case of the more general WKB approximation Hall (2013); Simon (2015); notably, a similar approach was applied to a generalization of our setting by Gaiotto et al. Gaiotto et al. (2011).
The claim of this first conjecture is somewhat weaker than that of (Beukers, 2007, Conj. 1). Namely, it does not make any claim that each lattice point in $\overline{\Lambda}$ must be associated to an eigenvalue of the Heun operator, or that such an eigenvalue (if it exists) must be unique; rather, it only claims that, if we know a priori that a Heun eigenvalue lies near a given lattice point, then it should be offset by the predicted asymptotic formula.
In general, this is all one should expect from the sort of asymptotic argument used to justify Conjecture V.1; the WKB approximation is not generally expected to converge to a precise answer, but it may provide an estimate of possible error in nearby solutions. One might then expect that, depending on the Heun equation, certain lattice points may have two or more associated eigenvalues, or none at all. On the other hand, by providing a rigorous proof in a particular case, Beukers also suggests that the following, stronger claim may be true:
Conjecture V.2.
There is a value $R>0$, possibly depending on the parameters $(m_{0},m_{1},m_{2},m_{3})$, such that for any $\ell_{0}\in\frac{2\pi}{\Delta}\overline{\Lambda}$ with $|\ell_{0}|>R$, a unique accessory parameter $B^{\prime}$ exists such that $|B^{\prime}-\ell_{0}^{2}|\leq\ell_{0}$ and such that the Darboux equation carries a single-valued, real-analytic solution. Moreover, $B^{\prime}$ satisfies the asymptotic estimate given by Conjecture V.1.
We highlight one subtlety with the two conjectures we have raised: Corollaries IV.1 and IV.2 appear to give three real algebraic constraints on our problem—corresponding to each of the traces $\operatorname{tr}(PQ)$, $\operatorname{tr}(QR)$, and $\operatorname{tr}(PR)$ being real—but the accessory parameter $B$ is defined by only two independent real parameters. In the case of the Lamé equation, this mismatch is handled naturally by the result of Beukers (Proposition II.4); if any two of the pairwise traces are real, the third follows. Thus, unitarity corresponds to two real algebraic constraints on two real parameters, and one expects a discrete set of solutions (as predicted in (Beukers, 2007, Conj. 1)).
We offer two pieces of evidence that a similar situation must hold in the general case, i.e., that Proposition III.3 (and thus Corollaries IV.1 and IV.2) can be strengthened to require only two pairwise traces to be real. First, we note that the argument underlying Conjecture V.2 relies on the equivalence between unitary monodromy and the existence of real-analytic solutions (Proposition IV.3), rather than on our trace conditions above. In turn, the WKB approach used to approximate real-analytic solutions in Conjecture V.2 gives rise to two real algebraic conditions at each order of approximation, just as in the Lamé case studied by Beukers (Beukers, 2007, Conj. 1).
Secondly, as we show shortly, numerical approximation offers strong, direct evidence of two facts:
  1. 1.
    The asymptotic values predicted in Conjecture V.2 correspond to unitary monodromy groups, despite the apparent mismatch of algebraic constraints discussed above.
  2. 2.
    If one ensures that the traces $\operatorname{tr}(P_{0}Q_{0})$ and $\operatorname{tr}(Q_{0}R_{0})$ are real in the setting of Corollary IV.2, then one consistently finds that the trace $\operatorname{tr}(P_{0}R_{0})$ is real within a small tolerance.
To investigate these claims, we leverage our analytical results to numerically approximate eigenvalues of the Heun equation. In short, Corollary IV.2 (and thus Theorem III.5 itself) transforms the accessory parameter problem (A’) into the following rootfinding problem:
$\operatorname{Im}(\operatorname{tr}(P_{0}Q_{0}))=\operatorname{Im}(\operatorname{tr}(P_{0}R_{0}))=\operatorname{Im}(\operatorname{tr}(R_{0}Q_{0}))=0,$ (23)
where $P_{0}$, $Q_{0}$, and $R_{0}$ are the rescaled local monodromy matrices defined in Corollary IV.2.
A naïve approach to this problem is to try to minimize the sum of squares of these three expressions. However, the resulting energy landscape does not lend itself well to traditional minimization schemes; it carries a multitude of nonzero local minima, for one, along with fine-scale structure in the vicinity of each true eigenvalue. While preconditioning may recover such an approach, it is outside the present scope.
Figure 1: The three contours about which we integrate the Heun equation (1), in order to recover its monodromy matrices. We begin at the common point $i=\sqrt{-1}$, travel to the right-most point of a radius-$1/5$ circle about each pole (fixing $a=-1$), and travel counter-clockwise around the circle before returning to our starting point.
Instead, we use a Newton-like approach to find simultaneous roots of two of these three expressions, following the proposal of Beukers for the Lamé equation Beukers (2007). Namely, given an initial estimate $B_{n}$ for the accessory parameter, we calculate matrices $P_{0}$, $Q_{0}$, and $R_{0}$ by integrating (1) numerically about each pole and rescaling the monodromy matrices appropriately. In our simulations below, we fix the third puncture at $a=-1$ and integrate about the contours shown in Figure 1, using a fourth-order Runge–Kutta method Butcher (2008) with step size $|dz|=4\times 10^{-4}$. With these matrices in hand, we define the traces $t_{PQ}(B_{n})=\operatorname{tr}(P_{0}Q_{0})$ and $t_{QR}(B_{n})=\operatorname{tr}(Q_{0}R_{0})$ and, given a small $h\in\mathbb{C}$, their approximate (holomorphic) derivatives
$\nu=\frac{t_{PQ}(B_{n}+h)-t_{PQ}(B_{n})}{h},\qquad\mu=\frac{t_{QR}(B_{n}+h)-t_{QR}(B_{n})}{h}.$
In our own simulations, we fix $h=10^{-5}\in\mathbb{R}$. To calculate our next estimate $B_{n+1}=B_{n}+\epsilon$, we want to ensure that $t_{PQ}(B_{n}+\epsilon)$ and $t_{QR}(B_{n}+\epsilon)$ are real. To first order, this criterion becomes
$\operatorname{Im}\left(t_{PQ}(B_{n})+\nu\epsilon\right)=0,\quad\operatorname{Im}\left(t_{QR}(B_{n})+\mu\epsilon\right)=0.$
Simultaneously solving both gives $\epsilon$ as
$\epsilon=\frac{\overline{\mu}\operatorname{Im}(t_{PQ}(B_{n}))-\overline{\nu}\operatorname{Im}(t_{QR}(B_{n}))}{\operatorname{Im}(\overline{\nu}\mu)}.$
Finally, we iterate $B$ as $B_{n+1}=B_{n}+\epsilon$ and repeat until convergence. If the third trace $t_{PR}(B)=\operatorname{tr}(P_{0}R_{0})$ is real within a given tolerance (here, $3\%$ relative error), we accept the converged value.
Figure 2: (a) Approximate square-roots of the eigenvalues of Lamé’s equation (5), found by starting our algorithm at each point $z_{\text{init.}}$ in the set (24) and iterating until convergence. All values satisfy (23) within a relative error of $1\%$. We show the lattice $\frac{2\pi}{\Delta}\overline{\Lambda}\approx 1.198\,\mathbb{Z}^{2}$ in black and several asymptotic values (provided by Proposition V.1) in red; these are exactly the values considered by Beukers Beukers (2007). In accordance with his conjecture, these eigenvalues appear to map one-to-one with lattice points; we zoom into a square of width $0.04$ about two lattice points, to demonstrate this close agreement. (b) A “convergence map” for the Lamé equation, showing the convergence properties of our algorithm in the (square root of the) region $[-50,50]\times[-50,50]\subset\mathbb{C}$. The position of each pixel represents the initial value of $B$ in our algorithm, and its color (using a standard HSV color transform) represents the final value after 20 iterations. We see that our algorithm breaks down along certain complex arguments as we move further away from the origin.
(a) $\gamma=\varepsilon=\delta=1/3$
(b) $\gamma=\varepsilon=\delta=2/3$
(c) $\gamma=\varepsilon=\delta=1$
Figure 3: Approximate square-roots of the eigenvalues of three Heun equations, all with $\alpha=\beta$. As parameters move away from the Lamé values ($\gamma=\delta=\varepsilon=1/2$), eigenvalues drift away from the lattice $\frac{2\pi}{\Delta}\Lambda$ in different directions.
We apply our algorithm first to the Lamé equation (i.e., $\gamma=\delta=\varepsilon=1/2$), where our algorithm reduces to that of Beukers Beukers (2007). To exhaustively find all eigenvalues in a large region, we start our algorithm independently at each point
$\sqrt{z_{\text{init.}}}/1.198\in\{1,2,3\}\times\{-1,0,1,2\}\subset\mathbb{C}$ (24)
and run it until convergence; the factor of $1.198$ was chosen to align with the leading order prediction $\frac{2\pi}{\Delta}\overline{\Lambda}\approx 1.198\,\mathbb{Z}^{2}$ of Proposition V.1. The square roots of our converged eigenvalues are shown Figure 2(a) in blue, along with the lattice $\frac{2\pi}{\Delta}\overline{\Lambda}$ in black and several asymptotic solutions in red, provided by Proposition V.1; the latter are exactly those of Beukers. We zoom into two example lattice points to demonstrate how the Lamé spectrum holds closely to the predicted, asymptotic values, consistent with both Conjecture V.1 and Conjecture V.2.
Figure 2(b) shows a “convergence map” of the same experiment, to clarify the global convergence properties of our algorithm. The position of each pixel represents an initial estimate of $\sqrt{B}$, and its color/brightness represents the converged value; here, we use the standard HSV domain coloring of the complex plane. The center of this figure is the origin, and we show both positive and negative square roots of $B$ for visual clarity. In accordance with Figure 2(a), the plane is largely broken up into large cells that converge to the same point, and these cells approximately form a regular lattice. We can see that the algorithm itself breaks down for certain complex arguments as we move far from the origin; the boundaries between different basins of convergence grow noisier and more blurred in these regions. We can see the effect of this on the right-hand side of Figure 2(b), where the algorithm fails to converge for certain larger eigenvalues.
Finally, Figure 3 shows results for Heun equations with $\alpha=\beta$ and with $\gamma,\delta,\varepsilon$ all at $1/3$, $2/3$, and $1$, respectively. Most eigenvalues appear to move only slowly away from their Lamé counterparts as parameters are changed. In particular, consistent with Conjecture V.2, we appear to maintain a close correspondence between eigenvalues and the lattice $\frac{2\pi}{\Delta}\overline{\Lambda}$. We also make the key observation that only the traces $t_{PQ}$ and $t_{QR}$ were used in identifying these values, and yet the third trace $t_{PR}$ always appears real within a $3\%$ relative error. As discussed above, these results provide strong evidence that Proposition III.3 and Corollaries IV.1 and IV.2 can be strengthened to require only two pairwise traces to be real, as Beukers identified in the case of the Lamé equation (Proposition II.4).
Acknowledgements.
We would like to recognize the role of the MIT PRIMES-USA program for connecting the three authors and facilitating this research. In particular, we would like to recognize Pavel Etingof (MIT) for his insights and helpful discussions.
Data Availability Statement
We have published Python implementations of our algorithms on a GitHub repository, as well as instructions for their use: https://github.com/ericc2023/HeunSimulation.